OpenStudy (anonymous):

Find the derivative with respect to x of the integral from 2 to x cubed of the natural log of x squared, dx.

OpenStudy (anonymous):

https://gyazo.com/d467d76cdbb0a105dfd1d1d0c7ea8024

OpenStudy (anonymous):

it is no wonder that mathematicians invented notation use the equation tool below

OpenStudy (anonymous):

I got \[3x^2[\ln(x^2)]\]

OpenStudy (anonymous):

Will wolfram be able to confirm this for me?

OpenStudy (anonymous):

that question is in serious error where did it i come from? the variable \(x\) is being used in two places to mean to entirely different things

OpenStudy (anonymous):

*two

OpenStudy (anonymous):

it should read \[\frac{d}{dx}\int _0^{x^3}\ln(t^2)dt\] tell you math teacher this is a big mistake on his/her part

OpenStudy (some.random.cool.kid):

wolfram is a great site for graphing however take @satellite73 advice instead because wolfram will not cover this. Plus I just tested it...

OpenStudy (anonymous):

so lets pretend that is what it is and solve it first off \(\ln(t^2)=2\ln(t)\) by a well known property of the log

OpenStudy (anonymous):

It is

OpenStudy (anonymous):

And I just saw my mistake.

OpenStudy (anonymous):

wolfram will not be able to do it because it is nonsense

OpenStudy (anonymous):

As soon as you posted that equation

OpenStudy (some.random.cool.kid):

give or take wolfram works 50% of the time for me but it will not help with this stuff i know that much.

OpenStudy (anonymous):

I got this \[3x^2[\ln(x^6)]\]

OpenStudy (anonymous):

in any case you need to know two facts: the derivative of the integral is the integrand, and the chain rule replace \(t\) by \(x^3\) and then multiply by the derivative of \(x^3 \)

OpenStudy (anonymous):

^^ I did

OpenStudy (anonymous):

yes, that is right, although i would have written \[18x^2\ln(x)\]

OpenStudy (some.random.cool.kid):

question? why 18? now im confused?

OpenStudy (anonymous):

Weird enough wolfram was able to do it originally but when I entered t instead of x it gave me nothing

OpenStudy (anonymous):

@Ephemera will explain that to you

OpenStudy (anonymous):

Oh it is 18 due to the properties of logarithmic functions

OpenStudy (some.random.cool.kid):

ah makes sense.

OpenStudy (anonymous):

:)